OpenStudy (anonymous):

what is the range of y=x^2-5 when the domain is {-1,0,3.5}

OpenStudy (anonymous):

We know that x^2 is always positive or 0 so this function will have its minimal value at x=0 where y=0^2-5=-5

OpenStudy (anonymous):

So the lowest possible value, the range starts from -5. At x=-1 y=(-1)^2-5=-4 At x=3.5 y=(3.5)^2-5=7.25

OpenStudy (anonymous):

So what is the solution?

OpenStudy (anonymous):

{-4,-5,7.25}

OpenStudy (anonymous):

Yes but I prefer to arrange the numbers in order: {-5,-4,7.25}